OpenStudy (anonymous):

Simplify the following expression. √-98 √-50

myininaya (myininaya):

\[i \sqrt{98} i \sqrt{50}=i^2 \sqrt{2*49} \sqrt{25*2}=(-1) \sqrt{49} \sqrt{25} \sqrt{2} \sqrt{2}\] \[=-(7)(5)(2)\]

myininaya (myininaya):

\[=-35(2)=-70\]